Find the sample proportion.

1. n = 80, x = 68

2. n = 1500, x = 785


1
Expert's answer
2022-05-25T09:10:29-0400

1.


"p=\dfrac{x}{n}=\dfrac{68}{80}=0.85"

2.


"p=\dfrac{x}{n}=\dfrac{785}{1500}=\dfrac{157}{300}"